# MegaRAG: Multimodal Knowledge Graph-Based Retrieval Augmented Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) enables large language models (LLMs) to dynamically access external information, which is powerful for answering questions over previously unseen documents. Nonetheless, they struggle with high-level conceptual understanding and holistic comprehension due to limited context windows, which constrain their ability to perform deep reasoning over long-form, domain-specific content such as full-length books. To solve this problem, knowledge graphs (KGs) have been leveraged to provide entity-centric structure and hierarchical summaries, offering more structured support for reasoning. However, existing KG-based RAG solutions remain restricted to text-only inputs and fail to leverage the complementary insights provided by other modalities such as vision. On the other hand, reasoning from visual documents requires textual, visual, and spatial cues into structured, hierarchical concepts. To address this issue, we introduce a multimodal knowledge graph-based RAG that enables cross-modal reasoning for better content understanding. Our method incorporates visual cues into the construction of knowledge graphs, the retrieval phase, and the answer generation process. Experimental results across both global and fine-grained question answering tasks show that our approach consistently outperforms existing RAG-based approaches on both textual and multimodal corpora. Our code is available on GitHub.
@@
Chi-Hsiang Hsiao1thanks: Equal contribution.   Yi-Cheng Wang [^1] footnotemark: 1   Tzung-Sheng Lin2 Yi-Ren Yeh3Chu-Song Chen1 1National Taiwan University  2E.SUN Financial Holding Co., Ltd. 3National Kaohsiung Normal University 1{r12922048, d13922033, chusong}@csie.ntu.edu.tw 2francis-17710@esunbank.com3yryeh@nknu.edu.tw
@@ 1 Introduction
Humans naturally integrate multiple modalities such as textual, visual, and layout to fluidly transition between abstract and detailed reasoning. However, multimodal large language models (MLLMs) bai2025qwen2; grattafiori2024llama; hurst2024gpt; team2023gemini, despite recent progress, remain limited by constrained context windows, restricting their ability to deeply process long-form, domain-specific content. E.g., interpreting a history textbook involves both conceptual insights and localized observations, which remains challenging for MLLMs.
@@
On the other hand, RAG can enhance LLMs by providing on-demand access to external knowledge. Early text-based RAG relied on sparse or dense retrieval but struggled with deep, multi-hop reasoning in multimodal documents. Recently, Graph-based RAG introduces structured abstraction via entity-relation graphs. With models like GraphRAG edge2024local and LightRAG guo2024lightrag, long-range knowledge retrieval of improved scalability are enhanced through KG-assisted retrieval pipelines. However, these methods excel in text-based multi-hop reasoning but remain constrained in handling complex, multimodal content. Current graph-based RAG methods face some key limitations. First, existing approaches remain unimodal, overlooking visual cues like diagrams, charts or maps, yielding disjointed representations that hinder multimodal reasoning. Additionally, due to context window constraints, most approaches segment documents into independent chunks, extracting entities separately rather than sequentially. This leads to fragmented KGs that miss cross-chunk relationships and key entities.
@@
To our knowledge, while recent studies have explored manually constructed multimodal knowledge graphs (KGs) for RAG-based question answering lee-etal-2024-multimodal, automatically building such KGs for RAG-assisted reasoning remains underexplored. To address this gap, we introduce MegaRAG, a multimodal, graph-based RAG method that enhances cross-modal reasoning.
@@
To better handle the association of different modalities in visual documents, more relations beyond text-to-texts need to be extracted, such as text-to-figures and figure-to-figure relations. Although the parallel-reading-then-combining strategy can refine entities and relations as in GraphRAG edge2024local and LightRAG guo2024lightrag, such refinement still relies on a single chunk while overlooking global document information. To address this limitation, we design a page-based, two-round approach for KG construction. Our solution initiates a KG by simply extracting entity-relation pairs in parallel for every page of a document using existing MLLMs, and the page-based relations are joined to form an initial graph. As the initial KG may not capture the inter-relationship between texts and visual elements sufficiently well, we conduct refinement processes in subsequent stage(s), where the initial KG(s) serve as global guidance to capture subtle relationships often lost in naïve, isolated extraction. In particular, to maintain scalability while incorporating long-range dependencies, we avoid injecting the entire initial KG into the MLLM inputs. Instead, we retrieve only a subgraph of the entire KG for each page, yielding a lightweight yet context-aware input. This strategy enables progressive improvement of the graph’s structural coherence, semantic coverage, and cross-modal grounding.
@@
Figure 1: Overview of our MegaRAG for MMKG construction and MMKG-augmented generation. (a) Initial Construction: Multimodal inputs from each page are processed by an MLLM to extract entities and relations $(E,R)^{0}_{i}$ in parallel. The page-level results are then joined by aligning identical entity names and relations, forming the initial document-level MMKG $\mathcal{G}^{0}$. (b) Refinement: Each page retrieves a subgraph $\mathcal{G}^{0}_{i}$ from $\mathcal{G}^{0}$ to assist the MLLM in refining the initial graph, yielding $\mathcal{G}^{1}$. (c) Indexing: The refined MMKG is encoded by an MMRAG’s retrieval approach into dense entity, relation, and page embeddings for efficient retrieval. (d) Retrieval & Answer Generation: A user query is parsed into low- and high-level keywords for retrieving relevant subgraphs and pages. These are fed into the MLLM for 2-stage answer generation.
@@ 2 Related Work
RAG with Raw Data Source. Early RAG methods guu2020retrieval; lewis2020retrieval retrieve text chunks from corpora to support answer generation, primarily relying on retrieval strategies either sparse or dense. Sparse methods exemplified by TF-IDF 10.1145/361219.361220 and BM25 10.1561/1500000019 depend on lexical heuristics to match queries with relevant text segments. They offer computational efficiency but lack deeper semantic comprehension. Dense techniques karpukhin2020dense; khattab2020colbert; santhanam2022colbertv2 project queries and documents into a shared embedding space, significantly improving retrieval performance of lexical variations. Subsequent works have enhanced this pipeline using LLM recently: HyDE gao-etal-2023-precise generates a hypothetical answer to enrich the retrieval query, Self-RAG asai2024selfrag introduces reflection tokens to enable adaptive retrieval and self-critique within a single LLM, while RQ-RAG chan2024rqrag decomposes the query into sub-queries to improve context coverage. Despite their strong performance on text-based RAG tasks, these methods often struggle with multimodal documents involving complex texts, layouts and visual elements.
@@
Multimodal RAG (MMRAG). To tackle the limitations, more recent studies have focused on multimodal retrieval methods that better retain the structural information of documents. DSE ma2024unifying treats document screenshots as unified inputs and directly encodes their visual layout, text, and images into a single vector embedding. ColPaLi faysse2024colpali continues this direction by encoding document images into multi-vector embeddings, effectively capturing fine-grained visual cues. Its variant, ColQwen, replaces the PaLI-Gemma beyer2024paligemma with Qwen2-VL wang2024qwen2 and achieves improved retrieval performance. Moving beyond retrieval, VisRAG yu2024visrag integrates MLLMs into the full RAG pipeline. Instead of extracting text, it embeds document images directly for retrieval and incorporates them into the generation stage, allowing the model to jointly reason over visual and textual content.
@@
The above methods excel in text-to-image retrieval but fail to solve tasks involving a mixture of single-modality (e.g., text-to-text), cross-modality (e.g., text-to-image), and fused-modality (text+image-to-text+image) retrieval. GME zhang2024gme tackles this by introducing a unified embedding model that encodes diverse modality combinations and enables flexible retrieval within a shared representation space.
@@
While these approaches significantly enhance document understanding, they neglect the long-range corpus-level structure, which is essential for handling complex, multi-hop QA tanaka2023slidevqa; yang2018hotpotqa.
@@
RAG with Knowledge Graph. Knowledge-augmented generation procko2024graph leverages KGs to provide structured, factual context for LLMs. Within this line of research, SubgraphRAG li2025simple enhances efficiency through lightweight scoring mechanisms for subgraph retrieval, while G-Retriever he2024g frames subgraph selection as a Steiner Tree optimization problem to support large-scale textual graphs. Gao et al. gao-etal-2022-graph employ a learning-to-rank approach to improve retrieval from KGs. While these methods advance graph-based retrieval, they depend on manually constructed KGs, which are costly to build and require substantial domain expertise. Moreover, static KGs are inherently limited in addressing queries that require corpus-level reasoning beyond fixed graph structures.
@@
To address this limitation, GraphRAG edge2024local proposes building KGs directly from raw text using LLMs, followed by a hierarchical community detection algorithm traag2019louvain to cluster semantically related nodes. During inference, it prompts the LLM to generate intermediate answers for each community summary, scores them by confidence, and aggregates the top responses into a final answer. Although this enables corpus-level reasoning, it incurs high computational cost due to repeated LLM queries over many community summaries. To improve efficiency, LightRAG guo2024lightrag introduces a two-stage retrieval process: it first extracts local and global keywords from the query, then retrieves relevant nodes and their surrounding subgraphs using dense retrieval. This design reduces the need for repeated LLM inference and significantly improves scalability. which introduces a hybrid RAG framework that alternates between naive and graph-based retrieval. TOG-2 ma2025thinkongraph introduces a hybrid RAG method that alternates between dense retrieval and graph reasoning. However, these approaches rely on manually curated KGs, which are costly to construct and limited in coverage.
@@
However, these KG-augmented RAGs rely solely on textual KGs, limiting their ability to handle multimodal content such as images. To overcome this limitation, multimodal knowledge graphs (MMKGs)10.1007/978-3-030-21348-0_30; zhangmultimodal enrich KGs by associating entities with aligned visual (e.g., images), numeric (e.g., dates, measurements), and textual descriptions. A representative benchmark10.1007/978-3-030-21348-0_30 introduces MMKGs that were constructed by linking overlapping entities via sameAS relations and annotating them with web-crawled images and numeric literals. MMKGs have demonstrated utility across tasks, including KG completion mousselly2018multimodal; xie2017image, recommendation systems sun2020multi, and image captioning zhao2023boosting.
@@
More recently, MMKGs have been integrated into RAG pipelines to support multimodal QA with LLMs. For instance, Lee et al. lee-etal-2024-multimodal utilized manually constructed MMKGs that encode visual and factual knowledge, enabling LLMs to reason over structured multimodal inputs. Although this study improves performance, it depends on manually built, domain-specific MMKGs that are costly to scale. No existing method using LLMs to construct MMKG for RAG, and current systems still struggle with open-ended reasoning beyond predefined graph structures. Building scalable, automatically constructed MMKGs that support open-domain, MMRAG remains a key challenge.
@@ 3.1 MMKG Construction
We define our MMKG as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes representing entities, and $\mathcal{E}$ is the set of edges denoting relations between entities. Given a document consisting of $N$ pages, we extract four types of content from each page $i$: text content $\mathrm{T}_{i}$, figure images $\mathrm{F}_{i}$, table images $\mathrm{B}_{i}$, and the full-page rendered image $\mathrm{I}_{i}$ (which captures the layout of the page). These elements are obtained using an off-the-shelf document analysis tool. We define the input for page $i$ as $\mathrm{P}_{i}=\{\mathrm{T}_{i},\mathrm{F}_{i},\mathrm{B}_{i},\mathrm{I}_{i}\}$, which serves as input to our graph construction pipeline.
@@ 3.2 Indexing and Retrieval
We adopt a unified retrieval framework that integrates graph structure, represented by entities and relations, along with page images within a shared embedding space to enable seamless cross-modal retrieval. Specifically, we use GME zhang2024gme, a multimodal encoder that jointly embeds textual and visual inputs. GME aligns all content types, including both textual and visual information, into a common vector space, supporting text-to-text and text-to-image retrieval through a unified representation.
@@
Graph Retrieval. To retrieve relevant knowledge, we adopt a dual-level retrieval strategy guo2024lightrag that targets both entities and relations. Given a user query, we first prompt the MLLM to extract two types of keywords: low-level keywords corresponding to specific entities, and high-level keywords that capture broader concepts. These keywords are then embedded by using the same GME model adopted during indexing. Both low-level and high-level keywords are combined into a single keyword list and used to query the entity vector store, retrieving the top-$k$ most relevant entities. In parallel, the top-$k$ most relevant relations, along with their associated source and target entities, are retrieved from the relation store. To further enrich the context, each retrieved entity is expanded by incorporating its one-hop neighbors from $\mathcal{G}^{1}$. The final set of entities and relations serves as input to the downstream reasoning module.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enables large language models (LLMs) to dynamically access external information, which is powerful for answering questions over previously unseen documents. Nonetheless, they struggle with high-level conceptual understanding and holistic comprehension due to limited context windows, which constrain their ability to perform deep reasoning over long-form, domain-specific content such as full-length books. To solve this problem, knowledge graphs (KGs) have been leveraged to provide entity-centric structure and hierarchical summaries, offering more structured support for reasoning. However, existing KG-based RAG solutions remain restricted to text-only inputs and fail to leverage the complementary insights provided by other modalities such as vision. On the other hand, reasoning from visual documents requires textual, visual, and spatial cues into structured, hierarchical concepts. To address this issue, we introduce a multimodal knowledge graph-based RAG that enables cross-modal reasoning for better content understanding. Our method incorporates visual cues into the construction of knowledge graphs, the retrieval phase, and the answer generation process. Experimental results across both global and fine-grained question answering tasks show that our approach consistently outperforms existing approaches on both textual and multimodal benchmarks. Our code is available on https://github.com/AI-Application-and-Integration-Lab/MegaRAG.
@@
Chi-Hsiang Hsiao1,thanks: Equal contribution.   Yi-Cheng Wang1,11footnotemark: 1   Tzung-Sheng Lin4 Yi-Ren Yeh2Chu-Song Chen1,3 1Department of Computer Science and Information Engineering, National Taiwan University 2Department of Mathematics, National Kaohsiung Normal University 3FinTech Center, National Taiwan University 4E.SUN Financial Holding Co., Ltd. 1{r12922048, d13922033, chusong}@csie.ntu.edu.tw 4francis-17710@esunbank.com2yryeh@nknu.edu.tw
@@ 1 Introduction
Humans naturally integrate multiple modalities such as textual, visual, and layout to fluidly transition between abstract and detailed reasoning. However, multimodal large language models (MLLMs) Bai et al. (2025); Grattafiori et al. (2024); Hurst et al. (2024); Team et al. (2023), despite recent progress, remain limited by constrained context windows, restricting their ability to process long-form, domain-specific content. Interpreting a history textbook involves both global and localized observations, which remains challenging for MLLMs.
@@
On the other hand, RAG can enhance LLMs by providing on-demand access to external knowledge. Early text-based RAG relied on sparse or dense retrieval but struggled with deep, multi-hop reasoning in multimodal documents. Recently, Graph-based RAG introduces structured abstraction via entity-relation graphs. With models like GraphRAG Edge et al. (2024) and LightRAG Guo et al. (2025), long-range knowledge retrieval of improved scalability are enhanced through KG-assisted retrieval pipelines. However, these methods excel in text-based multi-hop reasoning but remain constrained in handling complex, multimodal content. Current graph-based RAG methods face some key limitations. First, existing approaches remain unimodal, overlooking visual cues like diagrams, charts or maps, yielding disjointed representations that hinder multimodal reasoning. Additionally, due to context window constraints, most approaches segment documents into independent chunks, extracting entities separately rather than sequentially. This leads to fragmented KGs that miss cross-chunk relationships and key entities.
@@
To our knowledge, while recent studies have explored manually constructed multimodal knowledge graphs (KGs) for RAG-based question answering Lee et al. (2024), automatically building such KGs for RAG-assisted reasoning remains underexplored. To address this gap, we introduce MegaRAG, a multimodal, graph-based RAG method that enhances cross-modal reasoning.
@@
To better handle the association of different modalities in visual documents, more relations beyond text-to-texts need to be extracted, such as text-to-figures and figure-to-figure relations. Although the parallel-reading-then-combining strategy can refine entities and relations as in GraphRAG Edge et al. (2024) and LightRAG Guo et al. (2025), such refinement still relies on a single chunk while overlooking global document information. To address this limitation, we design a page-based, two-round approach for KG construction. Our solution initiates a KG by simply extracting entity-relation pairs in parallel for every page of a document using existing MLLMs, and the page-based relations are joined to form an initial graph. As the initial KG may not capture the inter-relationship between texts and visual elements sufficiently well, we conduct refinement processes in subsequent stage(s), where the initial KG(s) serve as global guidance to capture subtle relationships often lost in naïve, isolated extraction. In particular, to maintain scalability while incorporating long-range dependencies, we avoid injecting the entire initial KG into the MLLM inputs. Instead, we retrieve only a subgraph of the entire KG for each page, yielding a lightweight yet context-aware input. This strategy enables progressive improvement of the graph’s structural coherence and cross-modal grounding.
@@
Figure 1: Overview of our MegaRAG for MMKG construction and MMKG-augmented generation. (a) Initial MMKG construction: an MLLM extracts page-level entities/relations $(E,R)^{0}_{i}$ in parallel and merges them into $\mathcal{G}^{0}$. (b) Refinement: each page retrieves a subgraph $\mathcal{G}^{0}_{i}$ to refine the graph, producing $\mathcal{G}^{1}$. (c) Indexing: $\mathcal{G}^{1}$ and page images are encoded into dense entity, relation, and page embeddings. (d) Retrieval & generation: low-/high-level keywords retrieve relevant subgraphs/pages, followed by two-stage answer generation.
@@ 2 Related Work
RAG with Raw Data Source. Early RAG methods Guu et al. (2020); Lewis et al. (2020) retrieve text chunks from corpora to support answer generation, primarily relying on retrieval strategies either sparse or dense. Sparse methods exemplified by TF-IDF Salton et al. (1975) and BM25 Robertson and Zaragoza (2009) depend on lexical heuristics to match queries with relevant text segments. They offer computational efficiency but lack deeper semantic comprehension. Dense techniques Karpukhin et al. (2020); Khattab and Zaharia (2020); Santhanam et al. (2022) project queries and documents into a shared embedding space, significantly improving retrieval performance of lexical variations. Subsequent works have enhanced this pipeline using LLM recently: HyDE Gao et al. (2023) generates a hypothetical answer to enrich the retrieval query, Self-RAG Asai et al. (2024) introduces reflection tokens to enable adaptive retrieval and self-critique within a single LLM, while RQ-RAG Chan et al. (2024) decomposes the query into sub-queries to improve context coverage. Despite their strong performance on text-based RAG tasks, these methods often struggle with multimodal documents involving complex texts, layouts and visual elements.
@@
Multimodal RAG (MMRAG). To tackle the limitations, more recent studies have focused on multimodal retrieval methods that better retain the structural information of documents. DSE Ma et al. (2024) treats document screenshots as unified inputs and directly encodes their visual layout, text, and images into a single vector embedding. ColPaLi Faysse et al. (2025) continues this direction by encoding document images into multi-vector embeddings, effectively capturing fine-grained visual cues. Its variant, ColQwen, replaces the PaLI-Gemma Beyer et al. (2024) with Qwen2-VL Wang et al. (2024b) and achieves improved retrieval performance. Moving beyond retrieval, VisRAG Yu et al. (2025) integrates MLLMs into the full RAG pipeline. Instead of extracting text, it embeds document images directly for retrieval and incorporates them into the generation stage, allowing the model to jointly reason over visual and textual content.
@@
The above methods excel in text-to-image retrieval but fail to solve tasks involving a mixture of single-modality (e.g., text-to-text), cross-modality (e.g., text-to-image), and fused-modality (text+image-to-text+image) retrieval. GME Zhang et al. (2025) tackles this by introducing a unified embedding model that encodes diverse modality combinations and enables flexible retrieval within a shared representation space.
@@
While these approaches significantly enhance document understanding, they neglect the long-range corpus-level structure, which is essential for handling complex, multi-hop QA Tanaka et al. (2023); Yang et al. (2018).
@@
RAG with Knowledge Graph. Knowledge-augmented generation Procko and Ochoa (2024) leverages KGs to provide structured, factual context for LLMs. Within this line of research, SubgraphRAG Li et al. (2025) enhances efficiency through lightweight scoring mechanisms for subgraph retrieval, while G-Retriever He et al. (2024) frames subgraph selection as a Steiner Tree optimization problem to support large-scale textual graphs. Gao et al. Gao et al. (2022) employ a learning-to-rank approach to improve retrieval from KGs. While these methods advance graph-based retrieval, they depend on manually constructed KGs, which are costly to build and require substantial domain expertise. Moreover, static KGs are inherently limited in addressing queries that require corpus-level reasoning beyond fixed KGs.
@@
To address this limitation, GraphRAG Edge et al. (2024) proposes building KGs directly from raw text using LLMs, followed by a hierarchical community detection algorithm Traag et al. (2019) to cluster semantically related nodes. During inference, it prompts the LLM to generate intermediate answers for each community summary, scores them by confidence, and aggregates the top responses into a final answer. Although this enables corpus-level reasoning, it incurs high computational cost due to repeated LLM queries over many community summaries. To improve efficiency, LightRAG Guo et al. (2025) introduces a two-stage retrieval process: it first extracts local and global keywords from the query, then retrieves relevant nodes and their surrounding subgraphs using dense retrieval. This design reduces the need for repeated LLM inference and significantly improves scalability. which introduces a hybrid RAG framework that alternates between naive and graph-based retrieval. TOG-2 Ma et al. (2025) introduces a hybrid RAG method that alternates between dense retrieval and graph reasoning. However, these approaches rely on manually curated KGs, which are costly to construct and limited in coverage.
@@
However, these KG-augmented RAGs rely solely on textual KGs, limiting their ability to handle multimodal content such as images. To overcome this limitation, multimodal knowledge graphs (MMKGs) Liu et al. (2019); Zhang et al. (2023) enrich KGs by associating entities with aligned visual (e.g., images), numeric (e.g., dates, measurements), and textual descriptions. A representative benchmark Liu et al. (2019) introduces MMKGs that were constructed by linking overlapping entities via sameAS relations and annotating them with web-crawled images and numeric literals. MMKGs have demonstrated utility across tasks, including KG completion Mousselly-Sergieh et al. (2018); Xie et al. (2017), recommendation systems Sun et al. (2020), and image captioning Zhao and Wu (2023).
@@
More recently, MMKGs have been integrated into RAG pipelines to support multimodal QA with LLMs. For instance, MR-MKG Lee et al. (2024) utilizes manually constructed MMKGs that encode visual and factual knowledge, enabling LLMs to reason over structured multimodal inputs. Although effective, this approach depends on manually built, domain-specific MMKGs that are costly to scale. Query-driven MMKG Bu et al. (2025) takes a step toward automatic MMKG construction by dynamically building query-specific local graphs for online multimodal reasoning. Nevertheless, such local graph construction is mainly designed for short-form, query-centered settings and does not explicitly model document-level structure or cross-page dependencies in long-form documents. Therefore, building scalable MMKGs that can be constructed automatically and support open-domain multimodal RAG over long, visually rich documents remains a key challenge.
@@ 3.1 MMKG Construction
We define our MMKG as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes representing entities, and $\mathcal{E}$ is the set of edges denoting relations between entities. Given a document consisting of $N$ pages, we extract four types of content from each page $i$: text content $\mathrm{T}_{i}$, figure images $\mathrm{F}_{i}$, table images $\mathrm{B}_{i}$, and the full-page rendered image $\mathrm{I}_{i}$ (which captures the layout of the page). These elements are obtained using a document analysis tool. We define the input for page $i$ as $\mathrm{P}_{i}=\{\mathrm{T}_{i},\mathrm{F}_{i},\mathrm{B}_{i},\mathrm{I}_{i}\}$, which serves as input to our graph construction pipeline.
@@ 3.2 Indexing and Retrieval
We adopt a unified retrieval framework that integrates graph structure, represented by entities and relations, along with page images within a shared embedding space to enable seamless cross-modal retrieval. Specifically, we use GME Zhang et al. (2025), a multimodal encoder that jointly embeds textual and visual inputs. GME aligns all content types, including both textual and visual information, into a common vector space, supporting text-to-text and text-to-image retrieval through a unified representation.
@@
Graph Retrieval. To retrieve relevant knowledge, we adopt a dual-level retrieval strategy Guo et al. (2025) that targets both entities and relations. Given a user query, we first prompt the MLLM to extract two types of keywords: low-level keywords corresponding to specific entities, and high-level keywords that capture broader concepts. These keywords are then embedded by using the same GME model adopted during indexing. Both low-level and high-level keywords are combined into a single keyword list and used to query the entity vector store, retrieving the top-$k$ most relevant entities. In parallel, the top-$k$ most relevant relations, along with their associated source and target entities, are retrieved from the relation store. To further enrich the context, each retrieved entity is expanded by incorporating its one-hop neighbors from $\mathcal{G}^{1}$. The final set of entities and relations serves as input to the downstream reasoning module.